You take a personal loan of 1 year of $100,000 from the bank at nominal interest rate of 6%, interest to be paid monthly, and principal to be paid at the end of the year. what is the effective annual interest rate?

a) 6%

b) 6.14%

c) 6.17%

d) 6.25%

c) 6.50%

please show workings
Answer #1

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.06/12)^12-1

      = 1.0617-1

        = 0.0617*100= 6.17%

Therefore, the effective annual rate is 6.17%.

Hence, the answer is option c.

In case of any query, kindly comment on the solution